The area of a circle is calculated as πr squared. A square sheet of metal has sides 2p. The circle fits perfectly into the square : The area between the square and the circle is?


1
Expert's answer
2022-02-21T16:21:16-0500

As the circle fits perfectly the square with side 2p, the radius of the circle should be equal p.

Therefore the area of the circle is "\pi p^2" . The area of the square is "(2 p)^2 = 4 p^2"


So the area between the square and the circle is "(4 - \pi) p^2"